if Lars' Land Racers wanted a unit market share of 24.0% with no change in their Average Selling Price Per Unit, what would their Dollar Market share? Assume that the total market size in dollars and units does not change.


1
Expert's answer
2022-02-25T10:47:14-0500

Unit market share= "\frac{Units sold by a Co}{Total Market units}"

Let units sold be X

Let total market sales be Y units

Average selling price is P and remains unchanged

24% "= \frac{X}{Y}\times 100"

Y= "\frac{100X}{24}"

Dollar Market share= "\frac{No. of Dollars earned by the product}{Total dollar sales in the market}"

= "\frac{XP}{\frac{100XP}{24}}"

"{XP}\times{\frac{24}{100XP}}= 0.24"